OpenStudy (anonymous):

find the zeros of the quadratic y=6x(x-1)

OpenStudy (anonymous):

@Nnesha @triciaal

OpenStudy (anonymous):

The zeroes of the quadratic function y are the values of x that make y = 0.

OpenStudy (anonymous):

So what values of x make \[6x(x-1)\]equal to zero?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Yes, 0 is one way to make y = 0

OpenStudy (anonymous):

When you plug in 0 into\[6x(x-1)\] you get\[6*0*(0-1)\] which is 0

OpenStudy (anonymous):

There is one other value of x that makes\[6x(x-1)\] equal to 0

OpenStudy (triciaal):

y=6x(x-1) when y = 0 6 =0 or 0=0 or x-1 = 0

OpenStudy (anonymous):

How am i supposed to be looking at the numbers in the parenthesis? can i foil 6x into (x-1) or is that not how this works?

OpenStudy (triciaal):

when you multiply by zero the result is zero so if y = 0 then one of the factors must be 0 what are the 3 factors 6 and x and (x-1)

OpenStudy (triciaal):

are you ok now?

OpenStudy (anonymous):

no XD

OpenStudy (anonymous):

so i have y=6x(x-1) Y=6*0*(0-1) 6*0=0 but 0-1=-1 so how does that make y=0?

OpenStudy (triciaal):

because you are multiplying 6*0 = 0 and then multiply by -1 = 0*-1 = 0 0 * anything is = 0

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

How would i find the next zero?

OpenStudy (triciaal):

6 = 0 or x = 0 or (x-1) = 0 we know 6 not = 0 you chose x = 0 next option (x-1) = 0 so x = 0 + 1 = 1

OpenStudy (anonymous):

wow o my gosh that makes sense! so the second zero is 1

OpenStudy (triciaal):

excellent!! just remember if you multiply and get zero then one of the factor = zero

OpenStudy (anonymous):

is it possible to have more than 2 zeros?

OpenStudy (triciaal):

not sure what you mean but if the product is zero look at the factors you can have many factors to multiply

OpenStudy (triciaal):

above for this problem you had 3 factors to multiply and give a product of zero

OpenStudy (anonymous):

like how we had 3 factors but only we were only able to use 2 to get answers... could you have more than 2 zeros? or would you need more factors?

OpenStudy (triciaal):

it's still one 0 the value for y set each factor to =0 to determine the possible values of x that will make y = 0

OpenStudy (triciaal):

for example if you had y = (x-2)x(x-1)=0 then (x-2)=0 or x = 0 or (x-1) =0

OpenStudy (anonymous):

ok c: thank you

OpenStudy (triciaal):

you are always welcome.